OpenStudy (anonymous):

find the sum of the series 1 - e + e^2/2! - e^3/3! + e^4/4! - ...

OpenStudy (anonymous):

\[\sum_{0}^{?\infty}(-e)^n/(n!)\]

OpenStudy (anonymous):

the top shouldnt have a question mark but that should look like something u can solve

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

I can write the sum out for you: \[\sum_{k=0}^{\infty}\frac{(-1)^n x^n}{n!}\] That looks close to e^x. But e^x is: \[\sum_{k=0}^{\infty}\frac{x^n}{n!}\] So, pull the n power out: \[\sum_{k=0}^{\infty}\frac{(-x)^n}{n!}=e^{-x}\]

OpenStudy (anonymous):

Of course, for your case it would be e^(-e)

OpenStudy (anonymous):

Sorry, I didn't realize someone answered already :P

OpenStudy (anonymous):

np it helps with the steps

OpenStudy (anonymous):

n yours explained more anyway

OpenStudy (anonymous):

Thank you :D